3. If someone was born January 1st in the year 1492 and received $10,000 every day of his/her life. Calculate how much the individual would be worth on January 1st in the year 2020? Assume he/she is still alive and no interest is paid over the years. Give the answer in numbers and words. [4 marks]

Need the solution with step to step.
Answer #1

It is enough to find the number of days the person lived, that is the number of days from 1492 January 1 to 2020 January 1. First of all we have to find the number of leap years and non-leap years.

First consider 15th century: 1492-1500. The leap years are 1492, 1496. So out of 9 years 2 are leap years and the other 7 are non leap years.

In 16th century, there were 25 leap years, 1504, 1508,...,1600,and 75 non leap years. In 17th century, as 1700 is not divisible by 400, there were 24 leap years and 76 non leap years.

18th century- 24 leap years, 76 non-leap years

19th century- 24 leap years, 76 non-leap years

20th century- 25 leap years, 75 non-leap years.

In 21st century, from 2001 to 2020 January 1, there were 4 leap years and 15 non-leap years.

Hence total number of leap years=2+25+24+24+24+25+4 =127.

Total number of non-leap years = 7+75+76+76+76+75+15= 400.

Total number of days lived from 1492 Jan 1 to 2020 Jan 1= 127*366+400*365+1 = 192483.

Hence the individual would be worth on January 1st in the year 2020= $1924830000.